OpenStudy (firejay5):

You push a friend sitting on a swing with a force of 210N. She accelerates at a rate of 4 m/s^2. What is the mass of your friend? What is your friend's weight?

OpenStudy (zale101):

I think you would use F=MA Force= Mass * Acceleration for this question

OpenStudy (firejay5):

@Zale101 A. M = F/A M = 210N/4m/s^2 M = 52.5kg or 53kg B. F = M * A F = 53kg * 4m/s^2 F = 212N

OpenStudy (firejay5):

@satellite73